OpenStudy (anonymous):

3x+1/2=x/5 Solve each equation for the variable

OpenStudy (anonymous):

\[3x+\frac{ 1 }{ 2 }= \frac{ x }{ 5 }\]

OpenStudy (anonymous):

multiply through by a common factor of 2 and 5 ie 10 then: 30x+5=2x

OpenStudy (anonymous):

can you show me step by step

OpenStudy (radar):

You can do this, think of it as something like this: